OpenStudy (anonymous):

How would you do this problem? Suppose that after you are loaded into a Ferris wheel car, the wheel begins turning at 4 rpm. The wheel has diameter 12 m and the bottom seat of the wheel is 1 m above the ground. express the height h of the seat above the ground as a function of time t seconds after it begins turning.

OpenStudy (anonymous):

This one is probably gonna be a sine/cos wave function, right? The cosine wave equation is y=acos(b(x−c))+d You're one meter of the ground, the height is always positive, and the highest point is 13 meters, so we need to take 1 + our amplitude, which is 12, so d = 7, the amplitude is 6, meaning the derivation from zero, so a = -6. Now comes the tricky part, where we have to find b and c. As far as we're concerned, the ride starts at 0 on the x axis, so we can say that c = 0. Now y=-6cos(b(x))+7 solvable with the given information. We know that every 15 seconds, one rotation happens. A normal period can be shown by the equation\[b \times \left( x - c\right) = 2\pi\] c is 0, so \[b \times 15 = 2\pi\] \[b = \frac{ \pi }{ 15 }\] The full equation is \[h = -6\cos(\frac{ 2\pi }{ 15 } \times t) + 7\] ... I think.

OpenStudy (anonymous):

thank you so much. this finally makes sense